OpenStudy (anonymous):

Solve the equation for the interval [0°, 360°): 4cos^2(theta)sin(theta)+2cos^2(theta)=0

OpenStudy (anonymous):

Notice that \[4\cos ^2(\theta )\sin (\theta )+2 \cos ^2(\theta )=2\cos ^2(\theta ) (2 \sin (\theta )+1) =0 \]

OpenStudy (anonymous):

So \[ \cos ^2(\theta )=0 \\or\\ \sin(\theta) =-\frac 12 \]

OpenStudy (anonymous):

\[ \cos ^2(\theta )=0\\ \theta =\frac \pi 2\\ \theta =3\frac \pi 2\\ \]

OpenStudy (anonymous):

\[ \sin(\theta) =-\frac 12 \\ \theta=\frac { 7 \pi}6\\ \theta= \frac{11 \pi }{6} \]

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

It's supposed to be a set of 5 intervals, and its in radians. However, I do understand your answer.